# Safe Multi-Agent Reinforcement Learning for Behavior-Based Cooperative Navigation

### v2

@@ I Introduction
Cooperative navigation of unmanned vehicles has gained considerable attention due to its applications in missions such as search and rescue [1], surveillance [2], and payload transfers [3]. As outlined in the literature [4], cooperative navigation can be achieved through various strategies, including leader-follower, virtual-structures, and behavior-based approaches. Behavior-based methods [5, 6] offer more flexibility, as the robots’ actions adapt to their own observations. The goal in behavior-based cooperative navigation is to maintain relative distances, avoid collisions, and jointly reach target locations.
@@ II Related Work
Cooperative Navigation Using RL Several studies applied RL to achieve cooperative navigation using different formulations. In [7] the authors used multi-agents proximal policy optimization[22] along with curriculum learning, reference targets, and relative positions of the robots with respect to each other. Additionally, [19] applied PPO for formation control of quadrotors, incorporating individual goals for each unit. The authors in [8] used imitation learning along with RL to formulate a distributed formation controller based on a leader-follower scheme. In [9] the authors used policy gradients along with a graph convolutional network (GCN) and reference targets for each robot to achieve static formations of drones. [23] and [24] achieved formation control of maritime unmanned surface vehicles using deep deterministic policy gradient (DDPG) and deep Q-networks, respectively, based on leader-follower approaches. Different from the previous works, we use a multi-agent extension of the soft-actor-critic (SAC), since the SAC showed improved sample efficiency when compared to other approaches, including on-policy methods [25]. Additionally, we only use reference targets for the centroid of the formation and information about only two neighbors for each robot. More importantly, we ensure collision-free training by using MPC-based safety filters, and we use the attention mechanism to account for the interaction between the robots.
@@
Safety in MARL: Safety is a crucial aspect in MARL, with ongoing developments addressing this concern. [26] utilized control barrier functions (CBF) [27] along with multi-agent DDPG (MADDPG) for collision-free navigation of two agents. [12] explored the use of attention modules and decentralized safety shields based on CBF to reduce collisions in autonomous vehicle scenarios. Similarly, [13] implemented a centralized neural network as a safety layer with MADDPG in cooperative navigation. Moreover, [28] proposed using a predictive shielding approach, along with MADDPG, to ensure the safety of multi-agents in cooperative navigation scenarios. [29] presented RL-based model predictive control to achieve leader-follower formation control of mobile robots while ensuring their safety. In contrast to the previous works, we focus on behavior-based navigation where collisions between agents during navigation is more probable to occur, since the agents are required to navigate close to each other and not just have a single instance where their paths intersect. Additionally, the previous studies focused on the task of having individual reference targets, which can be achieved by training a single RL policy and then deploy it to several robots as in [30, 31, 32]. Our task, on the other hand, necessitates using a MARL framework so that the robots interact with each other during the training to rely only on a reference location for the centroid. While the previous studies fell short of real-world experiments, we conducted extensive experiments with both trained and untrained policies on real robots to demonstrate the safety of our framework, and the performance of our behavior-based modification.
@@ III Problem Statement
We consider the task of safely training a team of mobile robots to move the centroid of their formation to a desired target location while avoiding collisions with static obstacles and neighboring robots, and aiming to maintain predefined distances between robots. We assume that each robot is equipped with a lidar sensor to perceive its surroundings, each robot can localize itself within the environment, and that there is no pre-knowledge about the environment or existing obstacles. To ensure the safety of the robots at all times, each robot has its own safety layer to avoid collisions with its neighbors and the obstacles. The robots do not communicate with each other, but navigate cooperatively based on information between their centroid and the reference target, and the relative distances. Each robot receives relative distances and angles only about its two neighbors even if the team consists of more robots, as illustrated in Fig. 2. Finally, each robot gets the distance and relative angle to the closest obstacle and the target location of the centroid. We do not assume that the global information is available for each robot. This enables the scalability of our approach when applied to more robots, without retraining the policy.
@@ IV-A Multi-Agent Reinforcement Learning (MARL):
For each agent we define the observation space as in Fig. 2. The observation includes 40 equiangular lidar readings, the relative distances and angles to its two neighbors, the relative distance and angle to the nearest obstacle, calculated from the lidar scan, the relative distance and angle to the goal of the centroid, the distance from the centroid of the formation to the goal , and the actions of the robot for the previous time step which is the pair (). For the action space, we control the linear and angular velocities and . The reward function for each robot is defined as:
where:
and are sparse rewards indicating if the centroid reached the target and if the robot has collided or hasn’t been moving (stuck) for several consecutive steps, respectively.
is a continuous reward for maintaining the formation, calculated as the negative of the error between the reference distance among the agents and the actual distance.
is a continuous reward for keeping safe distance to the closest obstacle, calculated as the negative distance to the closest obstacle.
is a continuous reward for guiding the agents to move their centroid to the desired goal location, calculated as the negative distance between the centroid and the goal.
@@ IV-B Attention-Based Critics:
To effectively capture relative information among agents, we implement an attention module (Fig. 3(b)) following the framework of [35], which computes attention based on queries, keys, and values. Each agent’s observation is encoded into embeddings by an encoder network. These embeddings are input to the and networks to generate keys and values, respectively, while the querying agent’s embedding is processed by the network. The computed attention weights are then used to enhance the encoded observations, which are concatenated with the agent-specific outputs from the attention module and input into the critics to compute Q-values.
@@ IV-C Model Predictive Safety Filter:
To ensure the safety of the agents, we implement distributed shields based on nonlinear model predictive control (NMPC). The MPC utilizes a mathematical model of the robot (1) to calculate the predicted states based on the current state and action. At each time step, the MPC solves an optimization problem (2) to find the optimal control sequence that satisfies a set of predefined constraints on the states and actions. Only the first action of the control sequence is applied while the rest of the sequence is discarded. Prior safety related studies have employed the MPC as a safety filter [36, 28, 37] due to its capability to handle systems with multiple states, controls, and constraints. To minimize the intervention of the MPC-safety filter and avoid disrupting the exploration of the RL agent, the mentioned studies align the MPC actions with those of the RL agent. In our work, we additionally, penalize the RL agent for deviating from the MPC safe actions, which in turn teaches the RL agent to not rely on the MPC-safety filter as we show in Sec.V-C.
IV-C1 Safety Filter Formulation:
@@ IV-C2 Prediction Model:
IV-C3 Optimal Control Problem:
@@
where represents the prediction horizon. The notation indicates predictions at time , assuming the current time is . The optimization terms in (2a) are structured in three main components. The first term measures the deviation between the RL agent’s proposed action (ActionRL in Fig. 3(a)) and the initial MPC action (ActionMPC), optimizing for minimal discrepancy. The Second term minimizes the magnitude of future control signals , promoting smoother transitions. Third, the distances and between the robot and its two neighbors, as well as to the nearest obstacle, are penalized to ensure safety. The exponential function applied to distances has proven effective in enhancing obstacle avoidance, as shown in previous research [38]. Weight matrices , , and are manually tuned to balance the contributions of the cost terms.
Constraints (2b–2e) ensure adherence to system dynamics and operational limits, defined as follows: (2b) sets the initial state, (2c) enforces the robot’s dynamic model, (2d) and (2e) maintain the actions and states within predefined bounds, where X and A denote the allowable sets of states and controls, respectively.
@@ V Experimental Evaluation
Goals Reached W/o Obs. | Goals Reached W/ Obs. | S-Shaped Path | ||||||||
Agent | Success ↑ | Colls ↓ | Tout ↓ | Success ↑ | Colls ↓ | Tout ↓ | Goals Reached ↑ | Time (sec.) ↓ | Colls ↓ | Avg form err (m) ↓ |
Ours | 99.5% | 0 | 0.5% | 96.2% | 0 | 3.8% | 100% | 108.93 14.8 | 0 | 0.392 0.17 |
MASAC [40] | 58% | 8.9% | 33.1% | 51.8% | 26.3% | 21.9% | 75.8% | 134.63 21.3 | 2.8 3.65 | 1.25 0.15 |
MADDPG [41] | 9.5% | 57.7% | 32.8% | 6.6% | 81.5% | 11.9% | 12.4% | 118.9 28.39 | 4.16 3.81 | 4.05 4.06 |
DMPC [38] | 13.9% | 0 | 86.1% | 6.9% | 0 | 93.1% | 13.4% | 162.7 15.53 | 0 | 0.127 0.184 |
Can we execute our method without the MPC? | ||||||||||
Ours without MPC | 98.6% | 1.4% | 0 | 96% | 2.5% | 1.5% | 93.7% | 112.35 16.7 | 0.45 1.48 | 0.42 0.23 |
Integrating the MPC Safety Layer with the Baselines | ||||||||||
MASAC [40] (With MPC) | 67.7% | 0 | 32.3% | 60.3% | 0 | 39.7% | 79.9% | 129.54 38.7 | 0 | 0.98 0.18 |
MADDPG [41] (With MPC) | 10.4% | 0 | 89.6 % | 7.6% | 0 | 92.4% | 13.05% | 164.9 41.02 | 0 | 3.56 3.67 |
@@ V-A Training With the MPC Filter
To evaluate the MPC’s impact on agent training, we simulated two environments commonly used in the safe MARL literature [12, 13, 26, 28, 29]. The first environment is an empty walled setup to teach the robots to move the centroid to the goal. The second environment includes randomly placed obstacles for collision avoidance learning. We train the policies in both environments with and without the MPC. We utilized the same network parameters to isolate the MPC filter’s effects. Additionally, we introduced a penalty in the RL policy for deviations from the MPC-safe actions. Comparative metrics include average formation error, number of goals reached per episode, and number of collisions. Each episode terminates once the maximum number of steps is reached, or one of the robots collides or is stuck.
@@ V-B Testing Against Baselines in Simulation
Reaching a Single Goal: In the first scenario, we tested the ability of the robots to reach the desired goals starting from different configurations with and without obstacles in the environment. At the start of each episode, the locations of the robots, obstacles, and goal location are randomized. The results are summarized in Table I. Our approach outperforms the baselines in terms of all three metrics.
@@
Reaching a Single Goal | ||||||
Agent | Ours | MASAC [40] | ||||
Configuration | Success ↑ | Time (sec.) ↓ | Colls ↓ | Success ↑ | Time (sec.)↓ | Colls ↓ |
In-formation | 100% | 27.58 7.82 | 0 | 100% | 44.7 15.4 | 1.0 0 |
3-corners | 100% | 22.3 0.37 | 0 | 100% | 35.01 16.6 | 1.0 0 |
Centerline | 100% | 17.71 0.43 | 0 | 100% | 18.6 0.21 | 1.5 0.7 |
Collinear | 100% | 15.94 1.6 | 0 | 100% | 15.55 3.5 | 0.5 0.7 |
Face-to-face | 100% | 45.7 2.2 | 0 | 100% | 22 6.5 | 2.5 0.7 |
@@ V-D Integrating the MPC Safety Layer with the Baselines:
Goals Reaching With Obstacles | S-Shaped Path | |||||||
Agent | Success ↑ | Time (sec.) ↓ | Colls ↓ | Touts↓ | Goals reached ↑ | Completion time (sec.) ↓ | Collisions ↓ | Avg form err (m) ↓ |
Ours | 100% | 31.43 5.56 | 0 | 0 | 100% | 143.2 5.7 | 0 | 0.34 0.016 |
MASAC [40] | 0% | - | 100% | 0 | 96.2% | 212.2 53.8 | 7.3 1.52 | 0.58 0.35 |
@@ V-E Real-World Experiments
Reaching a Single Goal: We tested five different starting configurations with a target location for the formation centroid, as shown in Fig. 6. Each configuration was tested twice for each policy, with results summarized in Table II. Our approach consistently achieved zero collisions across all configurations, including unseen scenarios, demonstrating superior generalization and safety. Additionally, our method reached the targets faster than MASAC, as it better coordinates robot movements.
@@
Three | 99% | 0 | 1% |
Four | 97% | 0 | 3% |
Five | 91% | 0 | 9% |
Six | 79% | 0 | 21% |
Seven | 76% | 0 | 24% |
Eight | 70% | 0 | 30% |
In unseen scenarios, our approach required more maneuvers to avoid collisions, unlike MASAC, where robots reached the targets but failed to avoid collisions. For example, in the face-to-face configuration, our robots rotated and moved sequentially to avoid collisions, whereas MASAC robots collided and moved while touching, which explains the less time taken (these results are shown in the video). All tested configurations were reachable by both methods. We further tested more configurations that were not reachable by the baseline in the video. Both approaches were tested in obstacle avoidance scenarios. Fig. 7 show two of these scenarios, and the results are shown in Table III.
S-Shaped Path Following: In the S-shaped path scenario, Fig. 7(a), we run the test three times for each approach and recorded the number of collisions, the average formation error over the whole path, number of timeouts, as well as the time taken to complete the path. The results are summarized in Table III. Our approach was able to successfully reach all targets without making any collisions, and in less times compared to the MASAC. Additionally, our approach had less average formation error compared to the MASAC over all the runs.

### v3

@@ I Introduction
Cooperative navigation of unmanned vehicles has gained considerable attention due to its applications in missions such as search and rescue [1], surveillance [2], and payload transfers [3]. As outlined in the literature [4], cooperative navigation can be achieved through various strategies, including leader-follower, virtual-structures, and behavior-based approaches. Behavior-based methods [5, 6] offer more flexibility, as the robots’ actions adapt to their own observations. The goal in behavior-based cooperative navigation is to maintain relative distances, avoid collisions, and reach target locations.
@@ II Related Work
Cooperative Navigation Using RL Several studies applied RL to achieve cooperative navigation using different formulations. In [7] the authors used multi-agents proximal policy optimization[22] along with curriculum learning, reference targets, and relative positions of the robots with respect to each other. Additionally, [19] applied PPO for formation control of quadrotors, incorporating individual goals for each unit. The authors in [8] used imitation learning along with RL to formulate a distributed formation controller based on a leader-follower scheme. In [9] the authors used policy gradients along with a graph convolutional network (GCN) and reference targets for each robot to achieve static formations of drones. [23] and [24] achieved formation control of maritime unmanned surface vehicles using deep deterministic policy gradient (DDPG) and deep Q-networks, respectively, based on leader-follower approaches. Moreover, [25] proposed a motion planner based on a multi-agent extension of the soft-actor-critic (SAC) to coordinate multiple robots to reach their respective goals. Different from the previous works, we only use reference targets for the centroid of the formation and information about only two neighbors for each robot. More importantly, we ensure collision-free training by using MPC-based safety filters, and we use the attention mechanism to account for the interaction between the robots.
@@
Safety in MARL: Safety is a crucial aspect in MARL, with ongoing developments addressing this concern. [26] utilized control barrier functions (CBF) [27] along with multi-agent DDPG (MADDPG) for collision-free navigation of two agents. [12] explored the use of attention modules and decentralized safety shields based on CBF to reduce collisions in autonomous vehicle scenarios. Similarly, [13] implemented a centralized neural network as a safety layer with MADDPG in cooperative navigation. Moreover, [28] proposed using a predictive shielding approach, along with MADDPG, to ensure the safety of multi-agents in cooperative navigation scenarios. [29] presented RL-based model predictive control to achieve leader-follower formation control of mobile robots while ensuring their safety. In contrast to the previous works, we focus on behavior-based navigation where collisions between agents during navigation are more likely to occur, since the agents are required to navigate close to each other and not just have a single instance where their paths intersect. Additionally, the previous studies focused on the task of having individual reference targets, which can be achieved by training a single RL policy and then deploy it to several robots as in [30, 31, 32]. Our task, on the other hand, necessitates using a MARL framework so that the robots interact with each other during the training to rely only on a reference location for the centroid. While the previous studies fell short of real-world experiments, we conducted extensive experiments with both trained and untrained policies on real robots to demonstrate the safety of our framework, and the performance of our behavior-based modification.
@@ III Problem Statement
We consider the task of safely training a team of mobile robots to move the centroid of their formation to a desired target location while avoiding collisions with static obstacles and neighboring robots, and aiming to maintain predefined distances between robots. We assume that each robot is equipped with a lidar sensor to perceive its surroundings, each robot can localize itself within the environment, and that there is no pre-knowledge about the environment or existing obstacles. To ensure the safety of the robots at all times, each robot has its own safety layer to avoid collisions with its neighbors and the obstacles. The robots do not communicate with each other, but navigate cooperatively based on information about their centroid and the reference target, and the relative distances. Each robot receives relative distances and angles only about its two neighbors even if the team consists of more robots, as well as the target location of the centroid, as illustrated in Fig. 2. Finally, each robot calculates the distance and relative angle to the closest obstacle from the lidar scan. We do not assume that the global information is available for each robot. This enables the scalability of our approach when applied to more robots, without retraining the policy.
@@ IV-A Multi-Agent Reinforcement Learning (MARL):
For each agent we define the observation space as in Fig. 2. The observation includes 40 equiangular lidar readings, the relative distances and angles to its two neighbors, the relative distance and angle to the goal of the centroid, the distance from the centroid of the formation to the goal , the relative distance and angle to the nearest obstacle (calculated from the lidar scan), the desired reference distance between the robots, and the actions of the robot for the previous time step which is the pair (). For the action space, we control the linear and angular velocities and .
The reward function for each robot is defined as:
where and are sparse rewards indicating whether the centroid reaches the target and whether the robot collides or remains stuck for several steps, respectively. is a reward for maintaining formation, computed as the negative error between the reference distances and actual distances among agents. is a reward for maintaining a safe distance from the closest obstacle, given by the negative obstacle distance. is a reward for guiding agents toward the goal, calculated as the negative distance between the centroid and the goal.
@@ IV-B Attention-Based Critics:
To effectively capture relative information among agents, we implement an attention module (Fig. 3(b)) following the framework of [36], which computes attention based on queries, keys, and values. Each agent’s observation is encoded into embeddings by an encoder network. These embeddings are input to the and networks to generate keys and values, respectively, while the querying agent’s embedding is processed by the network. The computed attention weights are then used to enhance the encoded observations, which are concatenated with the outputs from the attention module and input into the critics to compute Q-values.
@@ IV-C Model Predictive Safety Filter:
To ensure the safety of the agents, we implement distributed shields based on nonlinear model predictive control (NMPC). The MPC utilizes a mathematical model of the robot (1) to calculate the predicted states based on the current state and action. At each time step, the MPC solves an optimization problem (2) to find the optimal control sequence that satisfies a set of predefined constraints on the states and actions. Only the first action of the control sequence is applied while the rest of the sequence is discarded. Prior safety related studies have employed the MPC as a safety filter [37, 28, 38] due to its capability to handle systems with multiple states, controls, and constraints. To minimize the intervention of the MPC-safety filter and avoid disrupting the exploration of the RL agent, the mentioned studies align the MPC actions with those of the RL agent. In our work, we additionally, penalize the RL agent for deviating from the MPC safe actions, which in turn teaches the RL agent to not rely on the MPC-safety filter as we show in Sec.6.
IV-C1 Safety Filter Formulation
@@ IV-C2 Prediction Model
IV-C3 Optimal Control Problem
@@
where represents the prediction horizon. The notation indicates predictions at time , assuming the current time is . The optimization terms in (2a) are structured in three main components. The first term measures the deviation between the RL agent’s proposed action (ActionRL in Fig. 3(a)) and the initial MPC action (ActionMPC), optimizing for minimal discrepancy. The second term minimizes the magnitude of future control signals , promoting smoother transitions. Third, the distances and between the robot and its two neighbors, as well as to the nearest obstacle, are penalized to ensure safety. The exponential function applied to distances has proven effective in enhancing obstacle avoidance, as shown in previous research [39]. The weight matrices , , and are tuned such that has higher weights than to ensure that the first action matches the proposed action . has the highest weight to maximize the cost on getting close to obstacles.
Constraints (2b–2e) ensure adherence to system dynamics and operational limits, defined as follows: (2b) sets the initial state, (2c) enforces the robot’s dynamic model from Eq. (1), (2d) and (2e) maintain the actions and states within predefined bounds, where X and A denote the allowable sets of states and controls, respectively.
@@ V Experimental Evaluation
Training Details: For the attention module, we used four attention heads and a hidden dimension of 128. For our approach and the RL baselines, we employed MLPs with two hidden layers of size (256, 256) for both the critics and the actors, using ReLU activation functions. The entropy coefficient in SAC is optimized online. The learning rate for all modules is set to , with a batch size of 128 and a replay buffer size of . The observation space and reward are fixed for all RL approaches. In simulations, we used the Turtlebot3 (TB3) robots, while in the real world experiments we used two TB2 robots and one TB4 robot.
| Goals Reached W/o Obs. | Goals Reached W/ Obs. | S-Shaped Path | ||||||||||||
| Agent | Success ↑ | Colls ↓ | Tout ↓ | Success ↑ | Colls ↓ | Tout ↓ |
| Ours | 99.5% | 0 | 0.5% | 96.2% | 0 | 3.8% | 100% | 108.93 14.8 | 0 | 0.392 0.17 | ||||
| MASAC [25] | 58% | 8.9% | 33.1% | 51.8% | 26.3% | 21.9% | 75.8% | 134.63 21.3 | 2.8 3.65 | 1.25 0.15 | ||||
| MADDPG [41] | 9.5% | 57.7% | 32.8% | 6.6% | 81.5% | 11.9% | 12.4% | 118.9 28.39 | 4.16 3.81 | 4.05 4.06 | ||||
| DMPC [39] | 13.9% | 0 | 86.1% | 6.9% | 0 | 93.1% | 13.4% | 162.7 15.53 | 0 | 0.127 0.184 | ||||
| Can we execute our method without the MPC? | ||||||||||||||
98.6% | 1.4% | 0 | 96% | 2.5% | 1.5% | 93.7% | 112.35 16.7 | 0.45 1.48 | 0.42 0.23 | ||||
| Integrating the MPC Safety Layer with the Baselines | ||||||||||||||
67.7% | 0 | 32.3% | 60.3% | 0 | 39.7% | 79.9% | 129.54 38.7 | 0 | 0.98 0.18 | ||||
10.4% | 0 | 89.6 % | 7.6% | 0 | 92.4% | 13.05% | 164.9 41.02 | 0 | 3.56 3.67 | ||||
@@ V-A Training With the MPC Filter
| No. of Obstacles | Success↑ | Timeouts ↓ |
| 4 | 90.3% | 9.7% |
| 5 | 86.7% | 13.3% |
| 6 | 83.8% | 16.2% |
| 7 | 82.3% | 17.7% |
| 8 | 71.4% | 28.6% |
| 3 Dynamic Obstacles | 81.4% | 18.6% |
| 4 Dynamic Obstacles | 71.3% | 28.7% |
To evaluate the MPC’s impact on agent training, we simulated two environments commonly used in the safe MARL literature [12, 13, 26, 28, 29]. The first environment is an empty walled setup to teach the robots to move the centroid to the goal. The second environment includes randomly placed cylindrical obstacles for collision avoidance learning. We train the policies in both environments with and without the MPC. We utilized the same network parameters to isolate the MPC filter’s effects. Additionally, we incorporated a penalty in the RL policy for deviations from MPC-safe actions. To maintain stable training despite the MPC’s modified actions, we store the RL action in the replay buffer and include the MPC action as the previous action in the observation space. Comparative metrics include average formation error, number of goals reached per episode, and number of collisions. Each episode terminates once the maximum number of steps is reached, or one of the robots collides or is stuck.
@@ V-B Testing Against Baselines in Simulation
Reaching a Single Goal: In the first scenario, we tested the ability of the robots to reach the desired goals starting from different configurations with and without obstacles in the environment. At the start of each episode, the locations of the robots, obstacles, and goal location are randomized. The results are summarized in Table I. Our approach outperforms the baselines in terms of all three metrics. Additionally, we evaluated our approach in more complex environments by introducing up to eight static obstacles, including cubes and wall sections, as well as dynamic obstacles, while training was limited to cylindrical obstacles. The robots avoided collisions in all cases, but timeouts increased with obstacle density, reflecting the system’s conservative behavior and the challenge of navigating constrained spaces. Results are in Table II.
@@ V-D Integrating the MPC Safety Layer with the Baselines:
| Reaching a Single Goal | ||||
| Agent | Ours | MASAC [25] | ||
| Configuration | Goals reached ↑ | Colls ↓ | Goals reached ↑ | Colls ↓ |
| In-formation | 100% | 0 | 100% | 1.0 0 |
| 3-corners | 100% | 0 | 100% | 1.0 0 |
| Centerline | 100% | 0 | 100% | 1.5 0.7 |
| Collinear | 100% | 0 | 100% | 0.5 0.7 |
| Face-to-face | 100% | 0 | 100% | 2.5 0.7 |
| Goals Reaching With Obstacles | S-Shaped Path | |||||||||||||||
Collisions ↓ |
| Ours | 100% | 0 | 0 | 0 | 100% | 143.2 5.7 | 0 | 0.34 0.016 | ||||||||
| MASAC [25] | 0% | 0 | 100% | 0 | 96.2% | 212.2 53.8 | 7.3 1.52 | 0.58 0.35 | ||||||||
@@ V-E Real-World Experiments
Reaching a Single Goal: We tested five different starting configurations with a target location for the formation centroid, as shown in Fig. 5. Each configuration was tested twice for each policy, with results summarized in Table III. Our approach consistently achieved zero collisions across all configurations, including unseen scenarios, demonstrating superior generalization and safety. Additionally, our method reached the targets faster than MASAC, as it better coordinates robot movements.
@@
| 3 | 99% | 0 | 1% |
| 4 | 97% | 0 | 3% |
| 5 | 91% | 0 | 9% |
| 6 | 79% | 0 | 21% |
| 7 | 76% | 0 | 24% |
| 8 | 70% | 0 | 30% |
In unseen scenarios, our approach required more maneuvers to avoid collisions, unlike MASAC, where robots reached the targets but failed to avoid collisions. For example, in the face-to-face configuration, our robots rotated and moved sequentially to avoid collisions, whereas MASAC robots collided and moved while touching, which explains the less time taken (these results are shown in the video). All tested configurations were reachable by both methods. We further tested more configurations that were not reachable by the baseline in the video. Both approaches were tested in obstacle avoidance scenarios. Fig. 7 show two of these scenarios, and the results are shown in Table IV.
S-Shaped Path Following: In the S-shaped path scenario, Fig. 7(a), we run the test three times for each approach and recorded the number of collisions, the average formation error over the whole path, number of timeouts, as well as the time taken to complete the path. The results are summarized in Table IV. Our approach was able to successfully reach all targets without making collisions, and in less time compared to the MASAC. Additionally, our approach had less average formation error compared to the MASAC over all the runs.